determine the resultant and direction of the following vectors: (apply the sine and cosine law). A=30 units 45degrees North of west, B=50 units 75 degrees south of east
1
Expert's answer
2020-12-21T11:00:51-0500

"R=\sqrt{30^2+50^2-2\cdot30\cdot50\cdot\cos5°}=20.28(units)" . Answer


"\frac{20.28}{\sin5°}=\frac{30}{sin\alpha}\to \alpha=7.4°"


"180\u00b0-5\u00b0-7.4\u00b0=167.6\u00b0"


So, "270\u00b0-45\u00b0-167.6\u00b0=57.4\u00b0" south of east. Answer